# ACORD: An Expert-Annotated Retrieval Dataset for Legal Contract Drafting

### v1

@@ Abstract
Information retrieval, specifically contract clause retrieval, is foundational to contract drafting because lawyers rarely draft contracts from scratch; instead, they locate and revise the most relevant precedent. We introduce the Atticus Clause Retrieval Dataset (ACORD), the first retrieval benchmark for contract drafting fully annotated by experts. ACORD focuses on complex contract clauses such as Limitation of Liability, Indemnification, Change of Control, and Most Favored Nation. It includes 114 queries and over 126,000 query-clause pairs, each ranked on a scale from 1 to 5 stars. The task is to find the most relevant precedent clauses to a query. The bi-encoder retriever paired with pointwise LLMs re-rankers shows promising results. However, substantial improvements are still needed to effectively manage the complex legal work typically undertaken by lawyers. As the first retrieval benchmark for contract drafting annotated by experts, ACORD can serve as a valuable IR benchmark for the NLP community.
@@ 1 Introduction
Contracts are the backbone of modern business, with millions created daily. A 2023 ALM and Bloomberg Law survey found that 43% of corporate counsels spend at least half their time on tasks like drafting, editing, and negotiating contracts (Bloomberg Law (2023); Pery (2019)). Lawyers rarely write contracts from scratch; they rely on finding and adapting relevant examples. High-quality precedents are crucial for handling complex clauses that involve nuanced legal concepts and require precise language. Mistakes in wording or missing key ideas can lead to disputes, liabilities, or invalid clauses.
Research shows that Large Language Models (LLMs) can effectively analyze legal contracts and identify issues within them (Hoffman and Arbel, 2023; Martin et al., 2024). However, LLMs struggle to draft contracts independently. Table 1 compares an LLM-generated clause with two top-rated precedent clauses from ACORD. Lawyer edits, shown with strike-through and underlined text, highlight two main issues: (1) the LLM uses “boilerplate” language that creates conflicts between clauses, and (2) it introduces uncommon phrasing not found in precedents.
@@
LLM-Generated Result ACORD 5-Star Precedent Clause ACORD 4-Star Precedent Clause
@@
c Attorney Note: liability cap should not be at purchase price paid in the event the buyer hasn’t paid in full or fails to pay timely. ACORD 5-star and 4-star precedent clauses fix this defect by capping the liability at the amount “paid or payable” or referring to the amount in the payment section of the contract.
@@
Table 1: Comparison between a clause generated by a large language model (LLM) and two precedent clauses rated 5-star and 4-star in the ACORD dataset. Deficiencies in the LLM-generated clause, such as inconsistent phrasing and omitted legal concepts, are detailed with lawyer annotations using strike-throughs and underlined text. Attorney notes in footnotes explain why the LLM-generated clause falls short.
@@
To enhance LLMs’ contract drafting abilities, we use Retrieval-Augmented Generation (RAG), which combines LLMs’ text generation with a retrieval system. RAG mimics how lawyers improve drafting by first searching for relevant precedents and then basing their drafts on those examples. This approach has been shown to reduce hallucinations and enhance performance (Lewis et al., 2021; Niu et al., 2024; Magesh et al., 2024). By referencing high-quality, tested precedents, LLMs can produce better contract drafts. We must evaluate how well models can retrieve the relevant contract clauses to ensure RAG can work effectively.
@@
Clause retrieval poses several challenges. Contracts are multi-layered, with sections and subsections that can span pages, often broken into paragraphs outlining obligations (what is required) and exceptions (what is not required). Cross-references within clauses add further complexity, such as defined terms (e.g., “as defined in Section 4”) or exceptions (e.g., “except as set forth in Sections 9 and 11(a)” or “notwithstanding anything to the contrary”). These references often point to sections located pages apart, yet they are crucial for understanding a clause’s meaning. Additionally, whether one precedent clause should rank higher than the others can be subjective and depends on the user’s individual experience, the industry they are in, and their unique needs, among other things. This subjectivity is reflected in the annotators’ disagreement rate of 21%. [^1] The fraction of ratings where the annotators significantly disagreed on the score or where they disagreed on the relevance.
@@
Datasets with expert-annotated, domain-specific benchmarks are essential for improving LLMs in clause retrieval and contract drafting. However, such datasets are rare due to the high cost of expert annotation and the confidentiality of proprietary contracts. Many existing legal contract benchmarks consist of multiple-choice Q&A formats, often derived from bar exam questions, and do not reflect the work lawyers perform, with none focusing specifically on contract drafting. To address this, we introduce ACORD, the Atticus Clause Retrieval Dataset, the first expert-annotated retrieval benchmark explicitly designed for contract drafting.
@@
ACORD addresses the challenge of retrieving complex and heavily negotiated contract clauses, such as Limitation of Liability, Indemnification, Most Favored Nation, and Termination for Convenience. These clauses are central to contract drafting and require careful negotiation and precise language. The task is to retrieve the most relevant precedent clauses for a given query, where the output is a ranked list of top-rated clauses from a large corpus. Figure 1 illustrates this retrieval process. The dataset was developed collaboratively by seasoned lawyers, student annotators, and machine learning researchers. It comprises 114 unique queries and over 126,000 query-clause pairs, each annotated with a 1-to-5-point relevance score by experts. This rigorous annotation ensures that ACORD is a robust and reliable benchmark for evaluating retrieval models in legal contexts.
@@ 2.1 Information Retrieval
Information retrieval (IR) is the process of locating relevant documents in response to a user query, as seen in web search tools such as Google and Bing. Historically, lexical-based retrieval has been foundational in IR (Hambarde and Proenca, 2023), with BM25 being one of the most popular ranking functions. BM25 scores document relevance based on term frequency and inverse document frequency while mitigating the impact of prevalent words (Robertson and Zaragoza, 2009). However, lexical-based retrieval methods face the “lexical gap”, where reliance on exact word matches often fails to capture relevant semantic connections such as synonymy and word-order-dependent relationships (Thakur et al., 2021). Recent approaches have moved toward neural IR to mitigate this, designed to capture semantic connections beyond simple lexical matches. These IR models fall into two primary categories:
  • Retriever Models: These models map queries and documents into a vector space, allowing pre-computed document representations to be indexed. For example, dense retrievers based on Transformer models (Vaswani et al., 2023) have shown robust performance in many tasks such as citation prediction, argument retrieval, summarization, duplicate detection, and question-answering (Thakur et al., 2021; Muennighoff et al., 2023). Another example is sparse embeddings that learned token-level contextualized representations using Transformers that still could be efficiently indexed using an efficient inverse index (Zhao et al., 2020; Formal et al., 2021).
  • Neural Reranking Models: These models enhance initial retrieval results (often from a BM25 or efficient retrieval model) by re-ordering the returned documents. Leveraging cross-attention mechanisms in models such as BERT has significantly boosted reranking performance (Devlin et al., 2019). ColBERT (Contextualized late interaction over BERT) further refines this approach by creating token-level embeddings for both queries and documents, using a maximum-similarity mechanism to identify relevant documents. However, this comes with higher computational costs (Devlin et al., 2019; Khattab and Zaharia, 2020). Recent research highlights that instruction-tuned large language models (LLMs) can surpass traditional supervised cross-encoders in zero-shot passage reranking tasks (Sun et al., 2023). In particular, listwise, pointwise, or pairwise reranking have received the most attention (Tang et al., 2024; Zhuang et al., 2024; Qin et al., 2024).
@@ 2.2 Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG)
Despite their advanced text generation capabilities, LLMs have not been widely adopted in contract drafting due to the risk of hallucination. Existing works (Wang et al., 2024; Dahl et al., 2024) show that the generations of LLMs can be unreliable, untrustworthy, and risky in many cases. Many recent studies focus on developing and enhancing various RAG techniques to reduce hallucination and improve model performance in text generation (Niu et al., 2024; Magesh et al., 2024). Given the absence of an information retrieval-specific dataset tailored to this task, the open question is whether these methodologies can effectively retrieve precedent clauses for contract drafting purposes.
@@ 2.4 Legal Contract Datasets
Prior datasets focus on clause extraction rather than retrieval. Chalkidis et al. (2017) & Leivaditi et al. (2020) provide datasets for extracting basic contract and lease information, not legal clauses, making them less useful for contract drafting. Hendrycks et al., 2021 introduces CUAD, a large-scale expert-annotated dataset with over 2,000 clauses across 41 categories from 510 commercial contracts. While useful for drafting, CUAD lacks ranking, an essential component of IR datasets. MAUD (Wang et al., 2023), ContractNLI (Koreeda and Manning, 2021), and LegalBench (Guha et al., 2023) focus on reading comprehension and reasoning, but would need extensive expert annotations for IR research. BigLaw Bench (Core, Workflow, and Retrieval) (Harvey AI, 2024) targets day-to-day legal tasks but is not specific to contract drafting and has an order of magnitude fewer annotations than ACORD.
ACORD seeks to provide a domain-specific expert-annotated large-scale IR benchmark for contract clause retrieval.
@@ 3 ACORD: A Contract Clause Retrieval Dataset
ACORD contains 114 unique queries across 9 clause categories. It has over 126,000 query-clause pairs, each rated with a 1-5 score. See Section C.3 for more details of the queries and expert-rated clauses.
@@
The labeling process follows five steps: (1) extraction, (2) retrieval, (3) scoring, (4) reconciliation, and (5) expansion.
First (Extraction), student annotators extract relevant clauses for the 9 categories from the Contract Corpus. Student annotators received 5-10 hours of training on the Limitation of Liability and Indemnification clauses. For the other 7 categories, we reused expert-annotated clauses from CUAD. Second (Retrieval), annotators search the Clause Corpus to find 10 relevant and 20 2-star irrelevant clauses per query. Some queries yielded only 5-8 relevant clauses. Third (Scoring), two expert attorneys and one student assign a 3- to 5-star relevance score to each query-clause pair based on a detailed four-page guideline. Irrelevant clauses mistakenly identified as relevant receive a 2-star rating from attorneys. Fourth (Reconciliation), a panel of 3-6 experienced attorneys reviews and reconciles scores for pairs with disagreements exceeding 2 points or disputes over relevance. The panel adjusts scores to stay within 2 points, and the final score is the average of the three annotators’ ratings. Annotator disagreement is tracked by the number of averaged scores revised during reconciliation. Lastly (Expansion), irrelevant 1-star clauses are added to each query as described in the “Query-Clause Pair Score” section above.
@@ 4.1 Setup
We use the following five standard information retrieval metrics to measure performance: NDCG@5, NDCG@10, 5-star precision@5, 4-star precision@5, 3-star precision@5. NDCG measures the models’ ranking quality by how well they rank the most relevant clauses at the top of the search results, focusing on the top results. The x-star-precision@5 metrics evaluate the precision within the top 5 results by counting how many of the top 5 results meet or exceed the respective relevance score. See Section F.1 for details and definitions of the performance metrics.
@@
We evaluate several stand-alone retriever methods, including BM25 (Robertson and Zaragoza, 2009), MiniLM bi-encoder (22.7M parameters) (Wang et al., 2020) and OpenAI text embedding (OpenAI, 2024b). BM25 is a computationally inexpensive lexical or keyword-based retriever. In BM25, documents have higher scores if they contain terms from the query. Terms that are rarer in the overall corpus have higher weighting, so they contribute more to the document score. MiniLM bi-encoder (66M parameters) is an embedding model trained via knowledge distillation from BERT on various embedding datasets, including MSMARCO.
We evaluate two-phase retrieval-reranker systems by reranking the top $100$ BM25 outputs with MiniLM cross-encoder (22.7M parameters), GPT-4o-mini, GPT-4o, Llama-3.2-1B (1.23B parameters), and Llama-3.2-3B (3.21B parameters) (OpenAI, 2024a; Meta, 2024) models.
@@
We finetune the MiniLM cross-encoder on the training dataset. We choose the learning rate and the number of updates via grid search on the validation NDCG@10 score. See Section F.2 for full training details.
@@ 4.2 Results
We show the results of most tested search systems on the test dataset in Table 3 with full results in Appendix D. Overall, a BM25 retriever paired with pointwise GPT-4o performs best, fine-tuned models perform better than models without fine-tuning, and embedding models perform better than lexical models. However, results vary significantly by query and model. Additionally, results using a pointwise reranking are generally much better than those using pairwise reranking in our experiments; see Table D-5 for a comparison.
@@
In retrieval-only systems, the large OpenAI text embedding model performed best. Large LLM rerankers paired with the bi-encoder retriever significantly outperformed other methods. The GPT4o reranker with bi-encoder-miniLM achieved the highest results, with an NDCG@5 of 79.1% and an NDCG@10 of 81.2%. The GPT4o reranker with BM25 was a close second, outperforming all other methods by a noticeable margin. Smaller LLMs performed much worse than larger ones; for instance, Llama 3.2 showed a difference of 40%-points in most metrics between the 1B and 3B models.
@@
We finetuned the MiniLM cross-encoder on the training data and observed a modest increase of 2.0% and 5.4% in NDCG@5 and NDCG@10 scores, respectively. Meanwhile, the 3- and 4-star precisions decrease slightly, with a 5.1% increase in 5-star precision. Refer to Figure 2 for results before and after fine-tuning, and Table D-2 in Section D.1 for results by category. We also fine-tune the cross-encoder when paired with the bi-encoder for the initial retrievals; the results are in Table D-4.
Figure 2: The performance of BM25 with the MiniLM cross-encoder before and after fine-tuning. We see improvements in NDCG and 5-star precision scores. There are minor declines in the 3-star and 4-star precision scores. These results underscore the importance of domain-specific fine-tuning for advancing clause retrieval capabilities; however, the improvements are not large, underlining that the problem is not solved. See Table D-2 in Section D.1 for results for each category.
@@
We measure the performance of the models across each of the 9 clause categories and under each query and find performance varies substantially. Some results are unexpected. For example, the query “indemnity of broad-based claims” achieved perfect scores for NDCG@5 and 3-star and 4-star precision@5 using BM25 paired with MiniLM cross-encoder, whereas “unilateral liability cap” and “as-is clause” got zero NDCG@5 score. This contradicts what we observed from human annotators: extracting “unilateral liability cap” and “as-is clauses” was easier than extracting relevant “indemnity of broad-based claims” clauses. Results for each clause category can be found in Section D.1.

### v2

# ACORD: An Expert-Annotated Dataset for Legal Contract Clause Retrieval

@@ Abstract
Contract clause retrieval is foundational to contract drafting because lawyers rarely draft contracts from scratch; instead, they locate and revise the most relevant precedent clauses. We introduce the Atticus Clause Retrieval Dataset (ACORD), the first expert-annotated benchmark specifically designed for contract clause retrieval to support contract drafting tasks. ACORD focuses on complex contract clauses such as Limitation of Liability, Indemnification, Change of Control, and Most Favored Nation. It includes 114 queries and over 126,000 query-clause pairs, each ranked on a scale from 1 to 5 stars. The task is to find the most relevant precedent clauses to a query. The bi-encoder retriever paired with pointwise LLMs re-rankers shows promising results. However, substantial improvements are still needed to manage the complex legal work typically undertaken by lawyers effectively. As the first expert-annotated benchmark for contract clause retrieval, ACORD can serve as a valuable IR benchmark for the NLP community.
@@ 1 Introduction
Contracts are the backbone of modern business, with millions created daily. A 2023 ALM and Bloomberg Law survey found that 43% of corporate counsels spend at least half their time on tasks like drafting, editing, and negotiating contracts (Bloomberg Law (2023); Pery (2019)). Clause retrieval is critical to these tasks because lawyers rarely write contracts from scratch; they rely on finding and adapting high-quality precedent clauses, especially when drafting complex clauses. Mistakes in wording or missing key ideas can lead to disputes, liabilities, or invalid clauses.
Research shows that Large Language Models (LLMs) can effectively analyze legal contracts and identify issues within them (Hoffman and Arbel, 2023; Martin et al., 2024; Choi et al., 2024; Schwarcz et al., 2025). However, LLMs struggle to draft contracts independently. Table 1 compares an LLM-generated clause with two top-rated precedent clauses from ACORD. Lawyer edits, shown with strike-through and underlined text, highlight two main issues: (1) the LLM-generated language creates conflicts between clauses, and (2) it introduces language not typically found in standard contracts.
@@
LLM-Generated Result ACORD 5-Star Clause ACORD 4-Star Clause
@@
c Attorney Note: liability cap should not be at purchase price paid in the event the buyer hasn’t paid in full or fails to pay timely. ACORD 5-star and 4-star clauses fix this defect by capping the liability at the amount “paid or payable” or referring to the amount in the payment section of the contract.
@@
Table 1: Comparison between an LLM-generated liability cap clause and two expert-rated clauses from ACORD. Attorney annotations (strike-through and underlined text) highlight deficiencies in the LLM output, demonstrating why retrieval of high-quality precedent clauses is essential for effective contract drafting.
@@
Retrieval-Augmented Generation (RAG), which combines LLMs’ text generation with a retrieval system, mimics how lawyers improve drafting by first searching for relevant precedent clauses to use as the base. This approach has been shown to reduce hallucinations and enhance performance (Lewis et al., 2021; Niu et al., 2024; Magesh et al., 2024). We seek to evaluate how well models retrieve relevant clauses to ensure RAG can work effectively.
@@
Clause retrieval poses several challenges. Contracts are multi-layered, with sections and subsections that can span pages, often broken into paragraphs outlining obligations (what is required) and exceptions (what is not required). Cross-references within clauses add further complexity, such as defined terms (e.g., “as defined in Section 4”) or exceptions (e.g., “except as set forth in Sections 9 and 11(a)” or “notwithstanding anything to the contrary”). These references often point to sections located pages apart, yet they are crucial for understanding a clause’s meaning. Additionally, whether one clause should rank higher than the others can be subjective and depends on the user’s individual experience, the industry they are in, and their unique needs, among other things. This subjectivity is reflected in the annotators’ disagreement rate of 21%. [^1] The fraction of ratings where the annotators significantly disagreed on the score or where they disagreed on the relevance. The value is common for hard datasets in the area. Under Labeling Process in Section 3, we explain how we mitigate this issue through annotator selections, guideline design, and expert-panel review.
@@
Datasets with expert-annotated, domain-specific benchmarks are essential for improving LLMs in clause retrieval and contract drafting. However, such datasets are rare due to the high cost of expert annotations and the confidentiality of proprietary contracts. Many existing legal contract benchmarks consist of multiple-choice Q&A formats, often derived from bar exam questions, and do not reflect the work lawyers perform, with none focusing specifically on contract drafting. To address this, we introduce ACORD, the Atticus Clause Retrieval Dataset, the first expert-annotated retrieval benchmark explicitly designed for contract drafting.
@@
ACORD addresses the challenge of retrieving complex and heavily negotiated contract clauses, such as Limitation of Liability, Indemnification, Most Favored Nation, and Termination for Convenience. These clauses are central to contract drafting and require careful negotiation and precise language. The task is to retrieve the most relevant precedent clauses for a given query, where the output is a ranked list of top-rated clauses from a large corpus. Figure 1 illustrates this retrieval process. The dataset was developed collaboratively by seasoned lawyers, student annotators, and machine learning researchers. It comprises 114 unique queries and over 126,000 query-clause pairs, each annotated with a 1-to-5-point relevance score by experts. This rigorous annotation ensures that ACORD is a robust and reliable benchmark for evaluating retrieval models in legal contexts. The estimated annotation cost of ACORD would be over US \$1 million when using the prevailing rates of \$550 per hour for attorneys and \$150 for non-attorneys.
@@ 2.1 Information Retrieval
Information retrieval (IR) is the process of locating relevant documents in response to a user query, as seen in web search tools such as Google and Bing. Historically, lexical-based retrieval has been foundational in IR (Hambarde and Proenca, 2023), with BM25 being one of the most popular ranking functions. BM25 scores document relevance based on term frequency and inverse document frequency while mitigating the impact of prevalent words (Robertson and Zaragoza, 2009). However, lexical-based retrieval methods face the “lexical gap”, where reliance on exact word matches often fails to capture relevant semantic connections such as synonymy and word-order-dependent relationships (Thakur et al., 2021). Recent approaches have moved toward neural IR, designed to capture semantic connections beyond simple lexical matches, to mitigate this. These IR models fall into two primary categories:
Retriever Models
Queries and documents are mapped into a vector space by the models, allowing pre-computed document representations to be indexed. For example, dense retrievers based on Transformer models (Vaswani et al., 2023) have shown robust performance in many tasks such as citation prediction, argument retrieval, and question-answering (Thakur et al., 2021; Muennighoff et al., 2023). Another example is sparse embeddings that learned token-level contextualized representations using Transformers that still could be efficiently indexed using an efficient inverse index (Zhao et al., 2020; Formal et al., 2021).
Neural Reranking Models
Initial retrieval results (often from e.g. a BM25 model) are enhanced by the models by reordering the returned documents. Leveraging cross-attention mechanisms in models such as BERT has significantly boosted reranking performance (Devlin et al., 2019). ColBERT (Contextualized late interaction over BERT) refines this approach by creating token-level embeddings for queries and documents, using a maximum-similarity mechanism to identify relevant documents. However, this comes with higher computational costs (Khattab and Zaharia, 2020). Recent research highlights that instruction-tuned LLMs can surpass traditional supervised cross-encoders in zero-shot passage reranking tasks (Sun et al., 2023). In particular, list-, point-, or pairwise reranking have received attention (Tang et al., 2024; Zhuang et al., 2024; Qin et al., 2024).
@@ 2.2 Large Language Models (LLMs) and Retrieval-Augmented Generation (RAG)
Despite their advanced text generation capabilities, LLMs have not been widely adopted in contract drafting due to the risk of hallucination. Existing works (Wang et al., 2024; Dahl et al., 2024) show that the generations of LLMs often can be unreliable, untrustworthy, and risky. Many recent studies focus on developing and enhancing various RAG techniques to reduce hallucination and improve model performance in text generation (Niu et al., 2024; Magesh et al., 2024; Schwarcz et al., 2025). Given the absence of an information retrieval-specific dataset tailored to this task, the open question is whether these methodologies can effectively retrieve precedent clauses for contract drafting purposes.
@@ 2.4 Legal Contract Datasets
Prior datasets focus on clause extraction rather than retrieval. Chalkidis et al. (2017) & Leivaditi et al. (2020) provide datasets for extracting basic contract and lease information, not legal clauses, making them less useful for contract drafting. Hendrycks et al., 2021 introduces CUAD, a large-scale expert-annotated dataset with over 2,000 clauses across 41 categories from 510 commercial contracts. While useful for drafting, CUAD is designed for clause classification and extraction and therefore lacks ranking, an essential component of IR datasets. MAUD (Wang et al., 2023), ContractNLI (Koreeda and Manning, 2021), and LegalBench (Guha et al., 2023) focus on reading comprehension and reasoning, but would need extensive expert annotations for IR research. BigLaw Bench (Core, Workflow, and Retrieval) (Harvey AI, 2024) targets day-to-day legal tasks but is not specific to contract drafting and has an order of magnitude fewer annotations than ACORD.
Aggarwal et al. (2021); Joshi et al. (2022); Lam et al. (2023); Alonso and Samara Chatzianastasiou (2024); Kasundra and Dhankhar (2024) have amongst others explored generating contract clauses by learning from data or applying constraints like logic rules or similar contract context. However, they generally rely on small datasets with simple queries and clauses. What is missing is a large-scale expert-rated benchmark that reflects how lawyers search for complex clauses, such as limitation of liabilities. ACORD fills this gap by providing the first large-scale benchmark for clause retrieval with queries crafted by experienced lawyers to fit their real needs.
@@ 3 ACORD: A Contract Clause Retrieval Dataset
ACORD contains 114 unique queries across the 9 clause categories. It has over 126,000 query-clause pairs, each rated with a 1-5 score. See Section C.3 for more details of the queries and expert-rated clauses.
@@
The annotation process followed five steps. See Section A.3 for additional details.
(1) Extraction: Student annotators, after receiving 5–10 hours of training, extracted relevant clauses for two contract categories from the Contract Corpus. For the remaining seven categories, we reused expert-annotated clauses from the CUAD dataset.
(2) Retrieval: Student annotators searched the Clause Corpus to retrieve relevant clauses for each query, aiming for 10 relevant (3- to 5-star) and 20 irrelevant (2-star) clauses per query. Some queries yielded only 5–8 relevant clauses.
(3) Scoring: Two experienced attorneys and one student annotator rated each query-clause pair using a detailed four-page rubric. Clauses mistakenly retrieved as relevant were assigned a 2-star rating.
(4) Reconciliation: For cases with rating discrepancies greater than 2 stars or disagreement on relevance, a panel of 3–6 experienced attorneys reviewed and adjusted scores to bring them within a 2-point range. Final scores were calculated as the average of the three individual ratings.
(5) Expansion: To augment the dataset, 1-star irrelevant clauses were added to each query using clauses from the CUAD dataset.
@@ 4.1 Setup
We use the following five standard information retrieval metrics to measure performance: NDCG@5, NDCG@10, 5-star precision@5, 4-star precision@5, 3-star precision@5. NDCG measures the models’ ranking quality by how well they rank the most relevant clauses at the top of the search results, focusing on the top results. The x-star-precision@5 metrics evaluate the precision within the top 5 results by counting how many of the top 5 results meet or exceed the respective relevance score. For NDCG, we change the scores from 1-5 to 0-4 to ensure the computed metrics correctly reflect the quality of returned results. See Section F.1 for details and definitions of the performance metrics.
@@
We evaluate several stand-alone retriever methods, including BM25 (Robertson and Zaragoza, 2009), MiniLM bi-encoder (66M parameters) (Wang et al., 2020) and OpenAI text embedding (OpenAI, 2024b). BM25 is a computationally inexpensive lexical or keyword-based retriever. Documents with terms from the query have higher scores; rarer terms in the overall corpus have higher weighting, so they contribute more to the document score. MiniLM bi-encoder is an embedding model trained via knowledge distillation from BERT on various embedding datasets, including MSMARCO.
We evaluate two-phase retrieval-reranker systems by reranking the top $100$ BM25222We focus on BM25 for these experiments as it is one of the most common retrievers. Due to space constraints, we move results for the MiniLM bi-encoder into Appendix D. outputs with the MiniLM cross-encoder (22.7M parameters), GPT-4o-mini, [^3] Using gpt-4o-mini-2024-07-18 GPT-4o, [^4] Using gpt-4o-2024-08-06 Llama-3.2-1B (1.23B parameters), and Llama-3.2-3B (3.21B parameters) (OpenAI, 2024a; Meta, 2024) models.
@@
The MiniLM cross-encoder is fine-tuned on the training dataset. We choose the learning rate and the number of updates via grid search on the validation NDCG@10 score. See Section F.2 for full training details.
@@ 4.2 Results
We show the results of most tested search systems on the test dataset in Table 3 with full results in Appendix D. Overall, a BM25 retriever paired with pointwise GPT-4o performs best, fine-tuned models perform better than models without fine-tuning, and embedding models perform better than lexical models. However, results vary significantly by query and model. Also, results using a pointwise instead of pairwise reranking are generally much better in our experiments; see Table D-5 for a comparison.
@@
In retrieval-only systems, the large OpenAI text embedding model performed best. BM25 or MiniLM bi-encoder with large LLM rerankers significantly outperformed other methods. MiniLM bi-encoder with GPT4o reranker achieved the highest results, with an NDCG@5 of 79.1% and an NDCG@10 of 81.2%, while BM25 with GPT4o reranker was a close second, outperforming all other methods by a noticeable margin. Smaller LLMs performed much worse than larger ones; Llama 3.2 showed a difference of 40%-points in most metrics between the 1B and 3B models.
@@
We finetuned the MiniLM cross-encoder on the training data and observed a modest increase of 2.0% and 5.4% in NDCG@5 and NDCG@10 scores, respectively. Meanwhile, the 3- and 4-star precisions decrease slightly, with a 5.1% increase in 5-star precision. See Figure 2 for results before and after fine-tuning, and Table D-2 in Section D.1 for results by category. We also fine-tune the cross-encoder when paired with the bi-encoder for the initial retrievals; the results are in Table D-4.
Figure 2: The performance of BM25 with the MiniLM cross-encoder before and after fine-tuning. We see improvements in NDCG and 5-star precision scores with minor declines in the 3-star and 4-star precision scores. These results underscore the importance of domain-specific fine-tuning for advancing clause retrieval capabilities; however, the improvements are not large, underlining that the problem is not solved. See Table D-2 in Section D.1 for results for each category.
@@
We measure the performance of the models across each of the 9 clause categories and under each query and find performance varies substantially. Some results are unexpected. For example, the query “indemnity of broad-based claims” achieved perfect scores for NDCG@5 and 3-star and 4-star precision@5 using BM25 paired with MiniLM cross-encoder, whereas “unilateral liability cap” and “as-is clause” got zero NDCG@5 score. This is contrary to what we observed from human annotators: extracting “unilateral liability cap” and “as-is clauses” was easier than extracting relevant “indemnity of broad-based claims” clauses. Results for each clause category can be found in Section D.1.